OpenStudy (clamin):

PLEASE HELP!!  The vertex of this parabola is at (-3, -1). When the y-value is 0, the x-value is 4. What is the coefficient of the squared term in the parabola's equation? a. 7 b. 3 c. -3 d. -7

OpenStudy (whpalmer4):

It appears there must have been an error copying the problem. None of those answers will provide the desired parabola.

OpenStudy (mathmale):

Notice how this problem does not state whether the parabola opens up or down, to the left or to the right. Plot the vertex and the given point and by doing so narrow down the possibilities. If it turns out that the parabola opens to the right, it's a horiz. parabola, and the appropriate equation is \[x-h=a(y-k)^2.\]

OpenStudy (mathmale):

Experiment. Substitute the coordinates of the vertex an.d those of the given point, and determine whether this is enough info to enable you to find the coefficient of the squared term. In the equation I've shared (above), that coeff. would be ' a '

OpenStudy (mathmale):

@clamin? Hello, whp!

OpenStudy (whpalmer4):

Ah, guilty of making an assumption that wasn't warranted! I take back my previous objection to the copying of the problem, the answer is there when you pick the right sort of parabola...

OpenStudy (whpalmer4):

thanks, @mathmale! I worked the problem several times in my head last night while going off to sleep, but it didn't occur to me to question my assumption that it opened up and down...

OpenStudy (clamin):

you are right the parabola opens to the right

Directrix (directrix):

Using this tip: the parabola opens to the right. (x-h) = a(y-k)^2 (h,k) = (-3,-1) x + 3 = a (y +1)^2 (4,0) is a point on the parabola 4 + 3 = a( 0 + 1)^2 7 = a*1 a = 7 See attached graph @clamin